# Ruelle-Pollicott Decay of Out-of-Time-Order Correlators in Many-Body Systems

### v2

@@ Abstract
The out-of-time-order correlator (OTOC) quantifies information scrambling in quantum systems and serves as a key diagnostic of quantum chaos. In one-body systems with a classical counterpart, the relaxation of the OTOC is governed by Ruelle-Pollicott resonances. For many-body systems lacking a semiclassical limit, recent studies have identified an analogous role played by the Liouvillian spectrum of weakly open extensions of the dynamics, where the slowest decay rate – the Liouvillian gap – encodes relaxation. Here we study the kicked Ising spin chain and show that the long-time exponential decay of the OTOC in the isolated system occurs at a rate equal to twice this intrinsic gap. This correspondence persists even in crossover regimes between integrability and chaos, demonstrating that the Liouvillian spectrum provides a unified framework for understanding relaxation and irreversibility in closed many-body quantum systems.
@@ I Introduction
In contrast, identifying the quantum analogue of RP resonances in generic many-body systems remains an open question. Several works [10, 11, 12] have introduced coarse-grained formulations of the operator space, revealing isolated eigenvalues inside the unit disk that can be interpreted as quantum RP resonances. While conceptually appealing, this approach becomes impractical for large systems due to the exponential growth of the operator space and the difficulty of defining an appropriate coarse-graining. A more recent development exploits translational invariance and quasi-momentum decomposition [13], enabling the identification of resonances within each symmetry sector.
@@
The magnitude of the leading RP resonance determines the approach to equilibrium of dynamical observables, including the out-of-time-order correlator (OTOC). The OTOC measures the spread of local perturbations and serves as a sensitive probe of information scrambling in quantum systems. Although the butterfly effect in quantum chaos is not universal [18], the relaxation of the OTOC at late times provides a robust signature of underlying dynamics. In one-body systems, the crossover from integrability to chaos can be characterized through the OTOC decay rate [19], a result that extends to systems with mixed classical dynamics [20].
Here we extend this connection to many-body systems. We study the kicked Ising spin chain, a paradigmatic model that exhibits both integrable and chaotic regimes depending on its parameters. Specifically, we analyze three complementary quantities: (i) the Liouvillian gap $\bar{g}$ extracted from a weakly dissipative version of the model, (ii) standard chaos indicators from spectral statistics and eigenstate delocalization, and (iii) the intermediate time exponential decay rate $\alpha$ of the OTOC in the isolated chain. We show that $\alpha\approx 2\bar{g}$ even in transitional regimes between integrability and chaos, establishing the Liouvillian gap as a qualitative (and quantitative) measure of integrability in quantum many-body systems.
@@ II Many-body System: Kicked Ising Spin Chain
To investigate the interplay between relaxation dynamics, quantum chaos, and dissipation, we consider the kicked Ising spin chain [21, 22], a paradigmatic model of many-body Floquet dynamics. This system is particularly well suited to our study because its parameters allow for a continuous interpolation between integrable and chaotic behavior.
@@
where the “free” and “kick” components take the form
@@
This Floquet representation conveniently describes both unitary and, when extended, weakly dissipative evolutions.
Throughout this work we set $J=h_{z}=\tau=1$ and vary the transverse field $h_{x}$, which controls the degree of chaoticity. These values ensure nontrivial dynamics while reducing the parameter space to a single effective control variable. All simulations are performed for system sizes up to $L=12$ spins.
@@ III Quantum Chaos Diagnostics
To characterize the degree of quantum chaos in the kicked Ising spin chain, we employ two complementary diagnostics: spectral statistics and eigenstate delocalization. These indicators enable us to distinguish between integrable, chaotic, and intermediate regimes, providing essential context for interpreting the behavior of the OTOC and its connection to the Liouvillian gap.
@@
These spectral and eigenstate-based chaos indicators provide a quantitative map of the dynamical regime of the system. We analyze their behavior as a function of the transverse field strength $h_{x}$, which controls the kick strength in the Floquet dynamics. To avoid spectral mixing between symmetry sectors, all analyses are performed within the even-parity subspace, ensuring a clearer characterization of the underlying dynamical behavior. This characterization serves as a reference for studying the Liouvillian gap and exploring potential connections between unitary and dissipative indicators of quantum chaos.
@@ IV Out-of-Time-Order Correlator
At intermediate times, the approach to saturation of $C(t)$, at (what we will call intermediate times) is reflected by an exponential decay of $O_{1}(t)$, which can be present even in systems which are not completely chaotic [20]. For one-body, completely chaotic systems, the decay is given by the classical RP resonances[19].
@@ V Liouvillian Gap
where $\gamma>0$ controls the dephasing strength and $\hat{H}(t)=\hat{H}(t+\tau)$ is the Floquet Hamiltonian.
@@
Figure 2: Liouvillian gap $g(\gamma)$ as a function of the dissipation strength $\gamma$ for system sizes $L=6$ (blue circles), $L=8$ (orange squares), and $L=10$ (green triangles) and $h_{x}=0.8168$. The dashed line indicate quadratic fit. For $L=10$, the extrapolated value of $\bar{g}$ is $0.1429$
@@
To extract $\bar{g}$ in the weak dissipation regime, we compute $g(\gamma)$ for several small values of $\gamma$ and perform a quadratic fit in $\gamma$. Following Eq. (16), the fit is applied to the largest available system size. The eigenvalues of the Floquet-Liouvillian spectrum, from which $g(\gamma)$ is obtained, are computed using the Arnoldi-Lindblad method described in Appendix A. This procedure provides the extrapolated value $\bar{g}\equiv\lim_{\gamma\to 0}\lim_{L\to\infty}g(\gamma)$, characterizing the intrinsic relaxation of the isolated chain.
@@
Parity-resolved analysis – We analyze the Liouvillian spectrum further within each parity subspace. Across the integrable-to-chaotic transition, the gaps of the two subspaces intersect when plotted as a function of the dissipation strength. As a result, the smallest nonzero eigenvalue – and hence the global Liouvillian gap – originates from the even subspace in certain parameter regions, while in others it comes from the odd subspace. This leads to an effective switching between sectors, so that the extracted global Liouvillian gap combines contributions from both subspaces depending on which hosts the slowest decaying mode. Importantly, this does not imply that the Arnoldi-Lindblad method is superior to the approach in Ref. [14]; rather, it allows us to probe parameter regimes – particularly the crossover from integrable to chaotic dynamics – where the behavior of the Liouvillian gaps can be resolved in detail.
@@
Figure 3 compares the Liouvillian gaps computed within each parity subspace – parity-even (blue empty circles) and parity-odd (orange empty squares) – with the global gap obtained via the Arnoldi-Lindblad method (green crosses). Solid lines serve as guides to the eye for the subspace gaps. As the dissipation parameter is varied, the dominant eigenvalue can switch between subspaces: in certain regions the even sector exhibits the smallest eigenvalue, while in others the odd sector becomes dominant. The Arnoldi method naturally captures this transition, since it always selects the eigenvalue with the smallest real part across all subspaces. Moreover, as the system size increases, the spectra of the two parity sectors become increasingly separated, making such crossings less probable. This behavior highlights that the Arnoldi approach consistently identifies the slowest decaying mode of the global dynamics, regardless of the underlying sectoral structure.
The case analyzed in Fig. 3(a) corresponds to the fully chaotic regime, where the Liouvillian gap curves for the two parity subspaces are clearly separated and no crossings occur, reflecting a well-defined hierarchy of decay rates between the sectors. In contrast, Fig. 3(b) shows an intermediate regime between integrability and chaos, where the curves partially overlap and crossings emerge, indicating that the slowest decaying mode can originate from either subspace depending on the dissipation strength. Both regimes are further illustrated in Fig. 4, highlighting how the Arnoldi-Lindblad method effectively captures the dominant relaxation dynamics and smoothly interpolates between different contributions as the system transitions from integrable to chaotic behavior.
@@ VI Relation between OTOC Decay, the Liouvillian Gap and Chaos Parameters
In this section, we bring together the chaos indicators, OTOC decay exponents, and the Liouvillian gap – into a single comparative framework. Previously, we have shown how to compute the normalized level – spacing ratio $\eta$ and participation ratio $\bar{\xi}_{E}$ across different transverse fields $h_{x}$, how to extract the intermediate time exponential decay exponent $\alpha$ of the local correlator $O_{1}^{zz}(l=1,t)$, and how to obtain the Liouvillian gap $\bar{g}$ via quadratic extrapolation in the weak – dissipation limit for $L=10$.
@@
Figure 4(a) displays the chaos indicators $\eta$ (blue upward triangles) and $\bar{\xi}_{E}$ (cyan diamonds) as functions of the transverse field $h_{x}$. Both quantities are computed from the unitary Floquet spectrum at system size $L=12$, and – owing to the symmetries of the model – are analyzed within the parity-even subspace defined with respect to the external reflection operator. These indicators clearly capture the crossover from integrability to chaos.
Figure 4(b) shows the comparison between the OTOC decay exponent $\alpha$ (red crosses) and twice the Liouvillian gap $2\bar{g}$ (orange empty circles) as functions of $h_{x}$. For each value of $h_{x}$, $\alpha$ is obtained by fitting $|O_{1}^{zz}(1,t)|$ to an exponential decay at intermediate times in the isolated chain of size $L=12$, while $\bar{g}$ is extrapolated from $g(\gamma)$ for $L=10$. The two quantities closely track each other across both fully chaotic and transitional regimes.
Overall, this figure highlights that the OTOC decay exponent remains approximately equal to twice the Liouvillian gap for all $h_{x}$ values where scrambling-induced relaxation persists. For $h_{x}>1.00$, the correspondence becomes less pronounced, likely due to finite-size effects – larger system sizes would be required to achieve quantitative convergence of both $\alpha$ and $\bar{g}$ in this regime. Nevertheless, even at these modest sizes, a clear qualitative and quantitative agreement is observed, underscoring the robustness of the relationship between the OTOC decay rate and the Liouvillian gap.
It is worth noting that the decay exponent $\alpha$ and the Liouvillian gap $\bar{g}$ qualitatively track the transition between dynamical regimes. A similar qualitative (and quantitative) behavior of the OTOC decay was previously reported for quantum maps in Ref. [20].
@@ VII Conclusions
We have investigated the connection between the decay of the out-of-time-order correlator (OTOC), the Liouvillian gap, and spectral indicators of chaos in many-body quantum systems, focusing on the kicked Ising spin chain. Our central finding is that the intermediate time exponential decay rate of the OTOC for generic, symmetry-unresolved operators equals twice the Liouvillian gap of the corresponding weakly open extension of the dynamics. This correspondence, previously established for systems with a semiclassical limit, is shown here to persist across the entire crossover from integrable to chaotic regimes, demonstrating that the Liouvillian spectrum provides a unifying description of relaxation and irreversibility in quantum many-body dynamics.
A detailed parity-resolved analysis confirms that the intrinsic Liouvillian gap extracted from the Arnoldi-Lindblad method reproduces the global relaxation rate of the system, even when the dominant decay mode shifts between symmetry sectors. This robustness highlights that both the OTOC and the Liouvillian gap respond coherently to changes in the underlying dynamical regime, providing consistent signatures of the transition from integrability to chaos.
@@
Overall, our results establish a quantitative bridge between unitary and dissipative descriptions of quantum relaxation, extending the Liouvillian-OTOC correspondence beyond the fully chaotic regime. This connection suggests that Liouvillian spectroscopy may serve as a general diagnostic tool for intermediate time dynamics and operator spreading in isolated quantum systems.

### v3

@@ Abstract
The out-of-time-order correlator (OTOC) quantifies information scrambling in quantum systems and serves as a key diagnostic of quantum chaos. In one-body systems with a classical counterpart, the relaxation of the OTOC is governed by Ruelle-Pollicott resonances. For many-body systems lacking a semiclassical limit, recent studies have identified an analogous role played by the Liouvillian spectrum of weakly open extensions of the dynamics, where the slowest decay rate – the Liouvillian gap – encodes relaxation. Here we study the kicked Ising spin chain and show that the long-time exponential decay of the OTOC in the isolated system occurs at a rate equal to twice this intrinsic gap. This correspondence is demonstrated across parameter regions exhibiting distinct level spacing statistics, indicating that the Liouvillian spectrum provides a robust framework for characterizing relaxation and irreversibility in closed many-body quantum systems.
@@ I Introduction
In contrast, identifying the quantum analog of RP resonances in generic many-body systems remains an open question. Several works [10, 11, 12] have introduced coarse-grained formulations of the operator space, revealing isolated eigenvalues inside the unit disk that can be interpreted as quantum RP resonances. While conceptually appealing, this approach becomes impractical for large systems due to the exponential growth of the operator space and the difficulty of defining an appropriate coarse-graining. A more recent development exploits translational invariance and quasi-momentum decomposition [13], enabling the identification of resonances within each symmetry sector.
@@
The magnitude of the leading RP resonance determines the approach to equilibrium of dynamical observables, including the out-of-time-order correlator (OTOC). The OTOC measures the spread of local perturbations and serves as a sensitive probe of information scrambling in quantum systems. Although the butterfly effect in quantum chaos is not universal [18], the relaxation of the OTOC at late times provides a robust signature of underlying dynamics. In one-body systems, the crossover from integrability to chaos can be characterized through the OTOC decay rate [19], a result that extends to systems with mixed classical dynamics [20]. More recently, a large-deviation theory of freeness has been proposed to describe the decay of higher-order OTOCs, introducing a hierarchy of characteristic time scales and mono- or multifractal behavior [21]. Within this framework, a possible interpretation of these behaviors in terms of Ruelle-Pollicott resonances has been suggested, but remains an open question. The connection was further generalized for higher-order OTOCS in minimal quantum circuits and Floquet systems, where exact results were obtained. In particular, for chaotic many-body models the OTOC decay rate is governed by the subleading eigenvalue of the corresponding quantum channel[22, 23].
Here we extend this connection to many-body systems. We study the kicked Ising spin chain, a paradigmatic model widely used in the study of quantum dynamics. Specifically, we analyze three complementary quantities: (i) the Liouvillian gap $\bar{g}$ extracted from a weakly dissipative extension of the dynamics, (ii) standard finite-size diagnostics based on spectral statistics and eigenstate delocalization, and (iii) the intermediate-time exponential decay rate $\alpha$ of the OTOC in the isolated chain. We show that $\alpha\approx 2\bar{g}$ across parameter regions displaying markedly different spectral and dynamical behavior. This establishes the Liouvillian gap as a robust indicator of long-time relaxation and operator spreading, independently of the specific finite-size statistical regime.
@@ II Many-body System: Kicked Ising Spin Chain
To investigate the interplay between relaxation dynamics, dissipation, and operator spreading in a many-body setting, we consider the kicked Ising spin chain [24, 25], a paradigmatic model of Floquet many-body dynamics. This system is particularly well suited to our study because its parameter space spans regions with markedly different spectral and dynamical properties, allowing us to test the robustness of our results across distinct regimes.
@@
where the free and kick components take the form
@@
This Floquet representation provides a convenient description of the unitary stroboscopic dynamics, which later serves as a natural starting point for incorporating weak dissipation at the level of the evolution superoperator.
Throughout this work we set $J=h_{z}=\tau=1$ and vary the transverse field $h_{x}$, which controls the degree of chaos. These values ensure nontrivial dynamics while reducing the parameter space to a single effective control variable. All simulations are performed for system sizes up to $L=12$ spins.
@@ III Quantum Chaos Diagnostics
To characterize the spectral and dynamical properties of the kicked Ising spin chain, we employ two complementary diagnostics: spectral statistics and eigenstate delocalization. These indicators are commonly used to assess chaotic versus integrable-like behavior in finite quantum systems and serve to organize the different parameter regions considered in our analysis, providing context for the behavior of the OTOC decay and its relation to the Liouvillian gap.
@@
These spectral and eigenstate-based diagnostics provide a quantitative characterization of the finite-size dynamical properties of the system. We analyze their behavior as a function of the transverse field strength $h_{x}$, which controls the kick strength in the Floquet dynamics. To avoid spectral mixing between symmetry sectors, all analyses are performed within the even-parity subspace, ensuring a consistent comparison across parameter values. This characterization provides a reference framework for the subsequent analysis of the Liouvillian gap and the OTOC decay.
@@ IV Out-of-Time-Order Correlator
At intermediate times, the approach of $C(t)$ to its saturation is well described by an exponential decay of $O_{1}(t)$, which can occur even in systems that are not fully chaotic [20]. For one-body systems with a well-defined classical limit, such as quantum maps, this decay has been shown to be governed by classical Ruelle-Pollicott resonances [19], which are extracted from a coarse-grained propagator acting on phase-space distributions.
In contrast, the kicked Ising spin chain considered here is a quantum many-body system that does not admit a classical phase-space description. As a consequence, the mechanisms controlling the long-time decay of the OTOC must be characterized using intrinsically quantum tools, which will be introduced in the following sections.
@@
The decay exponent $\alpha$ is extracted from the exponential relaxation of the OTOC in the intermediate-time window defined above. Importantly, this exponent is determined from a time regime in which the dynamics is no longer sensitive to the specific initial condition. In this window, transient effects have decayed while saturation has not yet set in, and the OTOC displays a well-defined exponential behavior.
This allows for a reliable and unambiguous extraction of the decay rate $\alpha$. Since the focus of this work is on the exponential relaxation preceding saturation, all reported values of $\alpha$ are consistently obtained from this intermediate-time regime.
@@ V Liouvillian Gap
where $\gamma>0$ controls the dephasing strength and $\hat{H}(t)=\hat{H}(t+\tau)$ is the time-dependent Hamiltonian.
@@
Figure 2: Liouvillian gap $g(\gamma)$ as a function of the dissipation strength $\gamma$ for system sizes $L=6$ (blue circles), $L=8$ (orange squares), and $L=10$ (green triangles), and $h_{x}=0.8168$. The dashed line indicates quadratic fit. For $L=10$, the extrapolated value of $\bar{g}$ is $0.1429$
@@
To extract $\bar{g}$ in the weak dissipation regime, we compute $g(\gamma)$ for several small values of $\gamma$ and perform a quadratic fit in $\gamma$. Following Eq. (16), the fit is applied to the largest available system size. The eigenvalues of the Floquet-Liouvillian spectrum, from which $g(\gamma)$ is obtained, are computed using the Arnoldi-Lindblad method described in the Appendix. This procedure provides the extrapolated value $\bar{g}\equiv\lim_{\gamma\to 0}\lim_{L\to\infty}g(\gamma)$, characterizing the intrinsic relaxation of the isolated chain.
@@
Parity-resolved analysis. We further analyze the Liouvillian spectrum by resolving it into the even and odd parity subspaces. As system parameters are varied, particularly across crossover regions between integrable-like and chaotic-like behavior, the leading gaps associated with these two subspaces may cross as functions of the dissipation strength. Consequently, the smallest nonzero Liouvillian eigenvalue—and therefore the global Liouvillian gap—can originate from either the even or the odd subspace, depending on the parameter regime. This results in an effective switching of the symmetry sector that hosts the slowest decaying mode when tracking only the leading relaxation rate. Importantly, this behavior reflects finite-size effects and does not signal any mixing or degeneracy between symmetry sectors. Rather, it indicates that different symmetry sectors can control the long-time relaxation in different parameter regions.
The Arnoldi-Lindblad approach employed here consistently captures the dominant Liouvillian eigenvalue governing relaxation, independently of the symmetry sector from which it originates. This allows us to resolve the behavior of the Liouvillian gap across a wide range of parameters, including finite-size crossover regimes where multiple symmetry sectors may compete in setting the slowest decay rate.
@@
Figure 3 compares the Liouvillian gaps computed within each parity subspace – parity-even (blue empty circles) and parity-odd (orange empty squares)-with the global gap obtained via the Arnoldi-Lindblad method (green crosses). Solid lines serve as guides to the eye for the subspace gaps. As the dissipation parameter is varied, the dominant eigenvalue can switch between subspaces: in certain regions the even sector exhibits the smallest eigenvalue, while in others the odd sector becomes dominant. The Arnoldi method naturally captures this transition, since it always selects the eigenvalue with the smallest real part across all subspaces. Moreover, as the system size increases, the spectra of the two parity sectors become increasingly separated, making such crossings less probable. This behavior highlights that the Arnoldi approach consistently identifies the slowest decaying mode of the global dynamics, regardless of the underlying sectoral structure.
The case analyzed in Fig. 3(a) corresponds to a parameter region where the Liouvillian gaps associated with the two parity subspaces are clearly separated over the range of dissipation strengths considered. In this regime, a well-defined hierarchy of decay rates emerges, and the slowest relaxation mode consistently belongs to a single symmetry sector. As a consequence, no crossings between the leading gaps of different parity sectors are observed. In contrast, Fig. 3(b) illustrates a parameter region displaying integrable-like finite-size behavior, where the Liouvillian gap curves associated with the two parity subspaces partially overlap and crossings occur as the dissipation strength is varied. In this case, the slowest decaying mode can originate from either subspace, leading to an effective switching of the sector that controls the long-time relaxation.
@@ VI Relation between OTOC decay, the Liouvillian gap and chaos parameters
In this section, we bring together the chaos indicators, OTOC decay exponents, and the Liouvillian gap - into a single comparative framework. Previously, we have shown how to compute the normalized level-spacing ratio $\eta$ and participation ratio $\bar{\xi}_{E}$ across different transverse fields $h_{x}$, how to extract the intermediate time exponential decay exponent $\alpha$ of the local correlator $O_{1}^{zz}(l=1,t)$, and how to obtain the Liouvillian gap $\bar{g}$ via quadratic extrapolation in the weak-dissipation limit for $L=10$.
@@
Figure 4(a) displays the chaos indicators $\eta$ (blue upward triangles) and $\bar{\xi}_{E}$ (cyan diamonds) as functions of the transverse field $h_{x}$. Both quantities are computed from the unitary Floquet spectrum at system size $L=12$, and—owing to the symmetries of the model—are analyzed within the parity-even subspace defined with respect to the external reflection operator. These indicators capture a crossover between integrable-like and chaotic-like spectral and eigenstate properties as the control parameter $h_{x}$ is varied.
Figure 4(b) shows the comparison between the OTOC decay exponent $\alpha$ (red crosses) and twice the Liouvillian gap $2\bar{g}$ (orange empty circles) as functions of $h_{x}$. For each value of $h_{x}$, $\alpha$ is obtained by fitting $|O_{1}^{zz}(1,t)|$ to an exponential decay at intermediate times in the isolated chain of size $L=12$, while $\bar{g}$ is extrapolated from $g(\gamma)$ for $L=10$.
The comparison shown in Fig. 4(b) is robust against the parity-resolved structure of the Liouvillian spectrum. As discussed above, depending on the parameter regime, the slowest decaying mode can originate from either the even or the odd parity subspace. The Arnoldi-Lindblad approach naturally captures this behavior by tracking the smallest nonzero Liouvillian eigenvalue, independently of the symmetry sector from which it arises. As a result, the extracted gap $\bar{g}$ consistently reflects the dominant relaxation rate that governs the long-time decay of the OTOC.
Standard spectral indicators of quantum chaos are included here as a complementary reference. While these quantities are known to be sensitive to finite-size effects, they provide an independent characterization of how the system parameters explored span regimes with qualitatively different finite-size dynamical behavior. Importantly, the agreement between the OTOC decay rate and twice the Liouvillian gap is found to persist across the entire parameter range, demonstrating the robustness of the Liouvillian-OTOC correspondence with respect to variations in the underlying finite-size dynamics.
Overall, this figure highlights that the OTOC decay exponent remains approximately equal to twice the Liouvillian gap for all $h_{x}$ values where scrambling-induced relaxation persists. For $h_{x}>1.00$, the correspondence becomes less pronounced, likely due to finite-size effects-larger system sizes would be required to achieve quantitative convergence of both $\alpha$ and $\bar{g}$ in this regime. Nevertheless, even at these modest sizes, a clear qualitative and quantitative agreement is observed, underscoring the robustness of the relationship between the OTOC decay rate and the Liouvillian gap.
It is worth noting that the decay exponent $\alpha$ and the Liouvillian gap $\bar{g}$ display a correlated dependence on the system parameters. A similar qualitative (and quantitative) behavior of the OTOC decay was previously reported for quantum maps in Ref. [20].
@@ VII Conclusions
In this work we have established a direct connection between the intermediate-time exponential decay of out-of-time-order correlators (OTOCs) and the spectral properties of the Liouvillian in a paradigmatic many-body quantum system, the kicked Ising spin chain. Our central finding is that the intermediate time exponential decay rate of the OTOC for generic, symmetry-unresolved operators equals twice the Liouvillian gap of the corresponding weakly open extension of the dynamics. While this relation was previously established in systems with a well-defined semiclassical limit, our results show that it also holds in a genuine many-body quantum system when tuning parameters across different dynamical regimes at finite system size, ranging from integrable-like to chaotic behavior. This shows that the Liouvillian spectrum provides a unified framework to characterize relaxation and irreversibility in quantum many-body dynamics.
A detailed parity-resolved analysis confirms that the intrinsic Liouvillian gap extracted from the Arnoldi-Lindblad method reproduces the global relaxation rate of the system, even when the dominant decay mode shifts between symmetry sectors. This robustness indicates that both the OTOC and the Liouvillian gap respond coherently to changes in the dynamical behavior of the system at finite size, providing consistent relaxation signatures across parameter regimes ranging from integrable-like to chaotic.
@@
Overall, our results establish a quantitative bridge between unitary and dissipative descriptions of quantum relaxation across a wide range of parameters. This connection suggests that Liouvillian spectroscopy may serve as a general diagnostic tool for intermediate-time dynamics and operator spreading in isolated quantum systems.